# Towards Quantum Accelerated Large-scale Topology Optimization

### v1

@@ Abstract
We present a new method that efficiently solves TO problems and provides a practical pathway to leverage quantum computing to exploit potential quantum advantages. This work targets on large-scale, multi-material TO challenges for three-dimensional (3D) continuum structures, beyond what have been addressed in prior studies. Central to this new method is the modified Dantzig-Wolfe (MDW) decomposition, which effectively mitigates the escalating computational cost associated with using classical Mixed-Integer Linear Programming (MILP) solvers to solve the master problems involved in TO, by decomposing the MILP into local and global sub-problems. Evaluated on 3D bridge designs, our classical implementation achieves comparable solution quality to state-of-the-art TO methods while reducing computation time by orders of magnitude. It also maintains low runtimes even in extreme cases where classical MILP solvers fail to converge, such as designs involving over 50 million variables. The computationally intensive local sub-problems, which are essentially Binary Integer Programming (BIP) problems, can potentially be accelerated by quantum computing via their equivalent Quadratic Unconstrained Binary Optimization (QUBO) formulations. Enabled by the MDW decomposition, the resulting QUBO formulation requires only sparse qubit connectivity and incurs a QUBO construction cost that scales linearly with problem size, potentially accelerating BIP sub-problem solutions by an additional order of magnitude. Due to limited access to quantum hardware, the validity of the formulated QUBO is verified using the Quantum Approximate Optimization Algorithm (QAOA). All observed and estimated speedups become increasingly significant with larger problem sizes and when moving from single-material to multi-material designs. This suggests that this new method, along with quantum computing, will play an increasingly valuable role in addressing the scale and complexity of real-world TO applications.
Keywords: Mixed-integer nonlinear programming; Mixed-integer linear programming; Binary integer programming; Quadratic unconstrained binary optimization; Topology optimization; Quantum computing
@@ 1 Introduction
As a PDE-constrained optimization problem, TO requires discretizing the design domain, typically via finite element methods (FEM), with each element representing a potential material site. This results in an optimization problem with a substantial number of binary (0/1) design variables, where 1 indicates material presence and 0 its absence. The nonlinear dependence of continuous state variables on these binary variables formulates TO as large-scale Mixed-Integer Nonlinear Programming (MINLP) problems, constrained by equality and inequality constraints. Being NP-hard in general [12], these problems are computationally intensive using traditional methods [13, 14, 15, 16, 17, 18, 19]. Our prior work introduced a TO framework [20] that significantly enhances computational efficiency by reducing optimization iterations and, consequently, the frequency of PDE solving. This framework integrates generalized Benders’ decomposition [21] with adaptive trust regions, formulating the master sub-problem (derived from the MINLP problem) as a multi-cut Mixed-Integer Linear Programming (MILP) problem. This approach enables estimation of the original objective function’s upper and lower bounds, accelerating solution convergence for both convex and non-convex problems across single- and multi-material designs. Compared to widely used traditional methods such as Solid Isotropic Material with Penalty (SIMP) [13], Floating Projection (FP) [17], and Sequential Approximate Integer Programming (SAIP) [18], this framework significantly reduces the number of optimization iterations and FEM analyses required, while achieving similar objective function values and optimized topologies. To scale this framework to larger TO problems, advancements to the MILP solver are required. Quantum computing, an emerging paradigm, offers potential superiority over classical computing for combinatorial optimization tasks [22, 23, 24, 25, 26, 27]. For example, quantum annealing can accelerate the search for global minima in quadratic unconstrained binary optimization (QUBO) problems by enabling faster escape from local optima [28, 29]. However, directly converting a MILP problem into an equivalent QUBO problem introduces new challenges. In TO, some design constraints often translate into global constraints within the MILP, which involve all design variables. When mapped into the QUBO formulation, these global constraints produce a dense Hamiltonian, i.e., a QUBO with a densely populated coefficient matrix. This not only imposes the need for all-to-all qubit connectivity, which exceeds the capabilities of most near-term quantum hardware [30, 31, 32], but also leads to high pre-processing costs. Specifically, constructing the QUBO’s dense coefficient matrix scales quadratically with the number of design variables, diminishing the overall efficiency and offsetting any potential quantum speedup.
@@
Therefore, this paper introduces a new method that significantly enhances the efficiency and scalability of solving MILP problems arising from large-scale TO. It also offers a practical pathway to harness quantum advantages by requiring only sparse qubit connectivity and ensuring that the cost of constructing QUBO scales linearly with problem size. This new method centers on a modified Dantzig-Wolfe (MDW) decomposition. Utilizing its block-angular structure, the MILP problem is decomposed into a series of reduced-size sub-problems, which are categorized as global or local, corresponding to global and local constraints, respectively. Unlike the standard Dantzig-Wolfe decomposition [33], the MDW formulation is designed to effectively control the size of the global sub-problem. The local sub-problems, which account for the majority of the computational cost, are essentially Binary Integer Programming (BIP) problems and can potentially be accelerated using quantum computing by translating them into equivalent QUBO problems. Importantly, each local sub-problem involves only local constraints, resulting in QUBO formulations with sparse Hamiltonians. This sparsity eliminates the need for all-to-all qubit connectivity, making the approach well-suited to near-term quantum hardware. Moreover, since each local sub-problem is independent, all QUBOs can be solved in parallel within each MDW iteration, enabling substantial gains in computational speed.
@@ 2 Preliminary
2.1 Formulation of Discrete Variable Topology Optimization (DVTO)
@@
$\nabla\cdot\bm{\sigma}(\mathbf{u})+\mathbf{b}=$ $\mathbf{0}\quad\text{in}\quad\Omega\;,$ (1)
$\bm{\sigma}(\mathbf{u})\cdot\mathbf{n}=$ $\mathbf{h}_{\Gamma}\quad\text{on}\quad\Gamma_{N}\;,$
$\mathbf{u}=$ $\mathbf{u}_{\Gamma}\quad\text{on}\quad\Gamma_{D}\;,$
@@
$\min_{\mathbf{u},\bm{\rho}}&\quad f(\mathbf{u},\bm{\rho})\\ \text{s.t.}&\quad\mathbf{K}(\bm{\rho})\mathbf{u}=\mathbf{f}\\ &\quad H_{i_{H}}(\bm{\rho})\leq 0,\quad i_{H}=1,\dots,n_{H}\\ &\quad\mathbf{u}\in\mathbb{R}^{3n_{v}},\bm{\rho}\in\{0,1\}^{n_{e}\times n_{M}}\end{aligned}\;,$ (2)
where $\mathbf{f}$ encompasses the discretized source terms and BCs; $\mathbf{u}=\{\mathbf{u}_{v}\}$ with $\mathbf{u}_{v}=\{u_{v},v_{v},w_{v}\}$ denoting the nodal displacement on each vertex of the FEM mesh; $\bm{\rho}=\{\rho_{e,m}\}$ with $\rho_{e,m}$ denoting the design variable defined at the centroid of each element $e$ in the mesh and for the candidate material $m$; $H_{i_{H}}$ denotes the inequality constraints that only exerted on design variables $\bm{\rho}$; and $\mathbf{K}$ is the linear operator discretized from Eq. (1). The mesh consists of $n_{v}$ nodal points (or vertices) and $n_{e}$ elements, with the design subject to $n_{M}$ candidate materials and $n_{H}$ inequality constraints.
@@
$\bm{\varepsilon}_{e}=$ $\left[\begin{matrix}\frac{\partial u_{v}}{\partial x}&\frac{\partial v_{v}}{\partial y}&\frac{\partial w_{v}}{\partial z}&\frac{\partial u_{v}}{\partial y}+\frac{\partial v_{v}}{\partial x}\end{matrix}\right.$ (3)
$\left.\begin{matrix}\frac{\partial u_{v}}{\partial z}+\frac{\partial w_{v}}{\partial x}&\frac{\partial v_{v}}{\partial z}+\frac{\partial w_{v}}{\partial y}\end{matrix}\right]^{\intercal}$
$=$ $\begin{bmatrix}\frac{\partial}{\partial x}&0&0&\frac{\partial}{\partial y}&\frac{\partial}{\partial z}&0\\ 0&\frac{\partial}{\partial y}&0&\frac{\partial}{\partial x}&0&\frac{\partial}{\partial z}\\ 0&0&\frac{\partial}{\partial z}&0&\frac{\partial}{\partial x}&\frac{\partial}{\partial y}\end{bmatrix}^{\intercal}\begin{bmatrix}u_{v}\\ v_{v}\\ w_{v}\end{bmatrix}$
$=$ $\mathbf{B}\mathbf{u}_{v}\;,$
@@
$\mathbb{C}_{m}=\frac{E_{m}}{(1+\nu_{m})(1-2\nu_{m})}\times$ (6)
$\begin{bmatrix}1-\nu_{m}&\nu_{m}&\nu_{m}&0&0&0\\ \nu_{m}&1-\nu_{m}&\nu_{m}&0&0&0\\ \nu_{m}&\nu_{m}&1-\nu_{m}&0&0&0\\ 0&0&0&\frac{1-2\nu_{m}}{2}&0&0\\ 0&0&0&0&\frac{1-2\nu_{m}}{2}&0\\ 0&0&0&0&0&\frac{1-2\nu_{m}}{2}\end{bmatrix}\;.$
@@
$&H_{M_{0}}(\bm{\rho})=\sum_{e=1}^{n_{e}}\sum_{m=1}^{n_{M}}\dfrac{\bar{M}_{m}}{n_{e}}\rho_{e,m}\leq\bar{M}_{\text{max}}\\ &H_{M_{e}}(\bm{\rho})=\sum_{m=1}^{n_{M}}\rho_{e,m}\leq 1,\quad e=1,2,\dots,n_{e}\end{aligned}\;,$ (8)
@@
The continuum TO problem in its discrete form as in Eq. (2) is essentially a mixed integer nonlinear programming (MINLP) problem. In our prior work [20], we developed a framework based on a multi-cut formulation and adaptive trust regions for efficiently solving it, which we refer to as DVTO-MT henceforth. This framework serves as the starting point for the methodology proposed in this work. Under this framework, at each iteration step (e.g., the $k$-th iteration) we solve a primal problem:
$\mathbf{u}^{k}=\mathbf{K}^{-1}(\bm{\rho}^{k})\mathbf{f}\;,$ (10)
@@
$t^{j}(\bm{\rho})=$ $\frac{1}{n_{M}n_{e}}\sum_{e=1}^{n_{e}}\left[\left(1-2\sum_{m=1}^{n_{M}}\rho_{e,m}^{j}\right)\left(\sum_{m=1}^{n_{M}}\rho_{e,m}\right)\right.$
$+$ $\left.\left(\sum_{m=1}^{n_{M}}\rho_{e,m}^{j}\right)^{2}\right]\leq d^{j}\;,$
@@
$\quad\bm{\rho}\in\{0,1\}^{n_{e}\times n_{M}}\;.$
[…]
@@
[…]
@@ 2.2 Sensitivity Analysis and Filtering
The gradient of the elemental stiffness matrix $\mathbf{K}_{e}$ with respect to the design variable $\rho_{e,m}$ is calculated as:
@@
[…]
@@
[…]

### v2

@@ Abstract
We present an efficient topology optimization (TO) method that not only enhances computational efficiency on classical computing but also provides a practical pathway for leveraging quantum computing to achieve further acceleration. The method targets large-scale, multi-material TO of three-dimensional (3D) continuum structures, beyond prior quantum TO studies limited to small-scale and single-material problems. Building on our discrete-variable TO framework (DVTO-MT), which employs multi-cut optimization and trust regions to reduce iteration counts and thereby PDE solver calls, the proposed method introduces a modified Dantzig–Wolfe (MDW) decomposition to further reduce per-iteration optimization time. The MDW method exploits the block-angular structure of the problem to decompose the mixed-integer linear program (MILP) into reduced-size global and local sub-problems. Evaluations on large-scale 3D bridge design problems demonstrate orders-of-magnitude reductions in computational time, with robust performance even for designs exceeding 50 million variables where classical MILP solvers fail to converge. Furthermore, the computationally intensive local sub-problems are transformed into equivalent quadratic unconstrained binary optimization (QUBO) formulations for quantum acceleration. The resulting QUBOs require only sparse qubit connectivity, a crucial consideration for near-term quantum hardware, and linear construction cost, offering the potential for an additional order-of-magnitude speedup. All observed and estimated speedups become more significant with increasing problem size and when extending from single-material to multi-material designs, highlighting the potential of the proposed method, coupled with quantum computing, to address the scale and complexity of real-world TO challenges.
keywords:
Topology optimization; Quantum computing; Mixed-integer programming; Combinatorial optimization; Quadratic unconstrained binary optimization
@@ 1 Introduction
As a PDE-constrained optimization problem, TO requires discretizing the design domain, typically via finite element methods (FEM), with each element representing a potential material site. This results in an optimization problem with a substantial number of binary (0/1) design variables, where 1 indicates material presence and 0 its absence. The nonlinear dependence of continuous state variables on these binary variables formulates TO as large-scale Mixed-Integer Nonlinear Programming (MINLP) problems, constrained by equality and inequality constraints. Being NP-hard in general [12], these problems are computationally intensive using traditional methods [13, 14, 15, 16, 17, 18, 19]. Our prior work introduced a TO framework [20] that significantly enhances computational efficiency by reducing optimization iterations and, consequently, the frequency of PDE solving. This framework, referred to as DVTO-MT, integrates generalized Benders’ decomposition [21] with adaptive trust regions, formulating the master sub-problem (derived from the MINLP problem) as a multi-cut Mixed-Integer Linear Programming (MILP) problem. This approach enables estimation of the original objective function’s upper and lower bounds, accelerating solution convergence for both convex and non-convex problems across single- and multi-material designs. Compared to widely used traditional methods such as Solid Isotropic Material with Penalty (SIMP) [13], Floating Projection (FP) [17], and Sequential Approximate Integer Programming (SAIP) [18], this framework significantly reduces the number of optimization iterations and FEM analyses required, while achieving similar objective function values and optimized topologies. To scale this framework to larger TO problems, advancements to the MILP solver are required. Quantum computing, an emerging paradigm, offers potential superiority over classical computing for combinatorial optimization tasks [22, 23, 24, 25, 26, 27]. For example, quantum annealing can accelerate the search for global minima in Quadratic Unconstrained Binary Optimization (QUBO) problems by enabling faster escape from local optima [28, 29]. However, directly converting a MILP problem into an equivalent QUBO problem introduces new challenges. In TO, some design constraints often translate into global constraints within the MILP, which involve all design variables. When mapped into the QUBO formulation, these global constraints produce a dense Hamiltonian, i.e., a QUBO with a densely populated coefficient matrix. This not only imposes the need for all-to-all qubit connectivity, which exceeds the capabilities of most near-term quantum hardware [30, 31, 32], but also leads to high pre-processing costs. Specifically, constructing the QUBO’s dense coefficient matrix scales quadratically with the number of design variables, diminishing the overall efficiency and offsetting any potential quantum speedup.
@@
Therefore, this paper introduces a new method that significantly enhances the efficiency and scalability of solving MILP problems arising from large-scale TO. It also offers a practical pathway to harness quantum advantages by requiring only sparse qubit connectivity and ensuring that the cost of constructing QUBO scales linearly with problem size. This new method centers on a modified Dantzig-Wolfe (MDW) decomposition. Utilizing its block-angular structure, the MILP problem is decomposed into a series of reduced-size sub-problems, which are categorized as global or local, corresponding to global and local constraints, respectively. Unlike the standard Dantzig-Wolfe decomposition [33], the MDW formulation is designed to effectively control the size of the global sub-problem. The DVTO-MT framework augmented with MDW can reduce computational time by orders of magnitude and maintains low runtimes even in extreme cases where classical MILP solvers fail to converge, such as designs with over 50 million variables. Furthermore, the local sub-problems, which account for the majority of the computational cost, are essentially Binary Integer Programming (BIP) problems and can potentially be accelerated using quantum computing by translating them into equivalent QUBO problems. Importantly, each local sub-problem involves only local constraints, resulting in QUBO formulations with sparse Hamiltonians. This sparsity eliminates the need for all-to-all qubit connectivity, making the approach well-suited to near-term quantum hardware. Moreover, since each local sub-problem is independent, all QUBOs can be solved in parallel within each MDW iteration, enabling substantial gains in computational speed.
@@
The remainder of this paper is organized as follows. Section 2 introduces the formulation of the DVTO-MT framework for multi-material TO of 3D continuum structures. Section 3 presents the proposed MDW methodology and its implementation for both classical and quantum computational paradigms. In Section 4, the proposed approach is evaluated and compared with existing methods in terms of solution quality and computational efficiency through large-scale 3D bridge design, including both single- and multi-material cases across various discretization resolutions. Finally, Section 5 concludes the paper by summarizing the study and highlighting its key contributions.
@@ 2 Preliminary
2.1 Formulation of DVTO-MT
@@
$\nabla\cdot\bm{\sigma}(\mathbf{u})+\mathbf{b}=\mathbf{0}$ $\text{in}\quad\Omega\;,$ (1)
$\bm{\sigma}(\mathbf{u})\cdot\mathbf{n}=\mathbf{h}_{\Gamma}$ $\text{on}\quad\Gamma_{N}\;,$
$\mathbf{u}=\mathbf{u}_{\Gamma}$ $\text{on}\quad\Gamma_{D}\;,$
@@
$\min_{\mathbf{u},\boldsymbol{\rho}}$ $\quad f(\mathbf{u},\bm{\rho})$ (2)
s.t. $\quad\mathbf{K}(\boldsymbol{\rho})\mathbf{u}=\mathbf{f}$
$\quad H_{i_{H}}(\bm{\rho})\leq 0,\quad i_{H}=1,\dots,n_{H}$
$\quad\mathbf{u}\in\mathbb{R}^{3n_{v}},\boldsymbol{\rho}\in\{0,1\}^{n_{e}\times n_{M}}$
where $\mathbf{f}$ encompasses the discretized source terms and BCs; $\mathbf{u}=\{\mathbf{u}_{v}\}$ with $\mathbf{u}_{v}=\{u_{v},v_{v},w_{v}\}$ denoting the nodal displacement on each vertex of the FEM mesh; $\boldsymbol{\rho}=\{\rho_{e,m}\}$ with $\rho_{e,m}$ denoting the design variable defined at the centroid of each element $e$ in the mesh and for the candidate material $m$; $H_{i_{H}}$ denotes the inequality constraints that only exerted on design variables $\bm{\rho}$; and $\mathbf{K}$ is the linear operator discretized from Eq. (1). The mesh consists of $n_{v}$ nodal points (or vertices) and $n_{e}$ elements, with the design subject to $n_{M}$ candidate materials and $n_{H}$ inequality constraints.
@@
$\bm{\varepsilon}_{e}=$ $\left[\begin{matrix}\frac{\partial u_{v}}{\partial x}&\frac{\partial v_{v}}{\partial y}&\frac{\partial w_{v}}{\partial z}&\frac{\partial u_{v}}{\partial y}+\frac{\partial v_{v}}{\partial x}&\frac{\partial u_{v}}{\partial z}+\frac{\partial w_{v}}{\partial x}&\frac{\partial v_{v}}{\partial z}+\frac{\partial w_{v}}{\partial y}\end{matrix}\right]^{\intercal}=\begin{bmatrix}\frac{\partial}{\partial x}&0&0&\frac{\partial}{\partial y}&\frac{\partial}{\partial z}&0\\ 0&\frac{\partial}{\partial y}&0&\frac{\partial}{\partial x}&0&\frac{\partial}{\partial z}\\ 0&0&\frac{\partial}{\partial z}&0&\frac{\partial}{\partial x}&\frac{\partial}{\partial y}\end{bmatrix}^{\intercal}\begin{bmatrix}u_{v}\\ v_{v}\\ w_{v}\end{bmatrix}=\mathbf{B}\mathbf{u}_{v}\;,$ (3)
@@
$\mathbb{C}_{m}=\frac{E_{m}}{(1+\nu_{m})(1-2\nu_{m})}\begin{bmatrix}1-\nu_{m}&\nu_{m}&\nu_{m}&0&0&0\\ \nu_{m}&1-\nu_{m}&\nu_{m}&0&0&0\\ \nu_{m}&\nu_{m}&1-\nu_{m}&0&0&0\\ 0&0&0&\frac{1-2\nu_{m}}{2}&0&0\\ 0&0&0&0&\frac{1-2\nu_{m}}{2}&0\\ 0&0&0&0&0&\frac{1-2\nu_{m}}{2}\end{bmatrix}\;$. (6)
@@
$H_{M_{0}}(\bm{\rho})=\sum_{e=1}^{n_{e}}\sum_{m=1}^{n_{M}}\dfrac{\bar{M}_{m}}{n_{e}}\rho_{e,m}\leq\bar{M}_{\text{max}}$ (8)
$H_{M_{e}}(\bm{\rho})=\sum_{m=1}^{n_{M}}\rho_{e,m}\leq 1,\quad e=1,2,\dots,n_{e}$
@@
The continuum TO problem in its discrete form as in Eq. (2) is essentially a MINLP problem. In our prior work [20], we developed the DVTO-MT framework based on a multi-cut formulation and adaptive trust regions for efficiently solving it. This framework serves as the starting point for the methodology proposed in the present work. Under this framework, at each iteration step (e.g., the $k$-th iteration) we solve a primal problem:
$\mathbf{u}^{k}=\mathbf{K}^{-1}(\boldsymbol{\rho}^{k})\mathbf{f}\;,$ (10)
@@
$t^{j}(\bm{\rho})=\frac{1}{n_{M}n_{e}}\sum_{e=1}^{n_{e}}\left[\left(1-2\sum_{m=1}^{n_{M}}\rho_{e,m}^{j}\right)\left(\sum_{m=1}^{n_{M}}\rho_{e,m}\right)+\left(\sum_{m=1}^{n_{M}}\rho_{e,m}^{j}\right)^{2}\right]\leq d^{j}\;,$
@@
$\quad\boldsymbol{\rho}\in\{0,1\}^{n_{e}\times n_{M}}\;.$
A branching procedure [20] automatically determines either (11) or (12) is solved for the master problem at each iteration step. The early optimization stage typically employs (12), whereas (11) is usually invoked during the later stages for finer topology adjustments.
@@
The iterative process of solving the primal problem (10) and the master problem (11) or (12) yields successive upper and lower bounds for the objective function in (2). To terminate the optimization, we minimize the gap between the lowest upper bound $U=\min_{j}(f(\mathbf{u}^{j},\bm{\rho}^{j}))$ for $j\leq k$ and the lower bound $\eta^{k}$ until it meets the condition $\tfrac{|\eta^{k}-U|}{|U|}<\xi$, with $\xi$ being the preset tolerance and set to $10^{-2}$ in this work.
@@ 2.2 Sensitivity Analysis and Filtering
The gradient of the elemental stiffness matrix $\mathbf{K}_{e}$ with respect to the design variable $\rho_{e,m}$ is calculated as [38, 39, 40]:
@@
$\begin{aligned} \dfrac{\partial f}{\partial\rho_{e,m}}=&\left(\dfrac{\partial f}{\partial\mathbf{u}}\right)^{\intercal}\cdot\dfrac{\partial\mathbf{u}}{\partial\rho_{e,m}}=\left(\dfrac{\partial f}{\partial\mathbf{u}}\right)^{\intercal}\dfrac{\partial\mathbf{K}^{-1}}{\partial\rho_{e,m}}\mathbf{f}=-\left(\dfrac{\partial f}{\partial\mathbf{u}}\right)^{\intercal}(\mathbf{K}^{-1})^{\intercal}\dfrac{\partial\mathbf{K}}{\partial\rho_{e,m}}\mathbf{K}^{-1}\mathbf{f}=-\left(\mathbf{K}^{-1}\dfrac{\partial f}{\partial\mathbf{u}}\right)^{\intercal}\cdot\dfrac{\partial\mathbf{K}}{\partial\rho_{e,m}}\mathbf{K}^{-1}\mathbf{f}\\ =&\bm{\mu}^{\intercal}\dfrac{\partial\mathbf{K}}{\partial\rho_{e,m}}\mathbf{u}=\bm{\mu}^{\intercal}\dfrac{\partial\mathbf{K}_{e}}{\partial\rho_{e,m}}\mathbf{u}=\bm{\mu}^{\intercal}\left(\int_{\Omega_{e}}\mathbf{B}^{\intercal}\mathbb{C}_{m}\mathbf{B}\mathrm{d}\Omega\right)\mathbf{u}\end{aligned}\;.$ (16)
@@
where $h_{e,e^{\prime}}(r)=\max(0,r-\|\mathbf{x}_{e}-\mathbf{x}_{e^{\prime}}\|_{2})$; $e^{\prime}$ denotes a neighbor element of $e$; and $r$ is the prescribed radius and set to $3h$ in this work with $h$ the mesh’s side length.